# Bias in Large Language Models: Origin, Evaluation, and Mitigation

### v1

@@ 2.1.2 Bias in Data Collection Methods
The intrinsic biases can also stem from the methods that are used to collect the training corpus. A large training corpus of LLMs often comprises of sub-datasets that are collected from different sources, including web scraping, social media and forums, books and literature, code repositories, scientific and academic papers, legal and finance documents, dialogue and transcriptions, government and public records, custom and proprietary database, and model-generated datasets. These different types of sources have their own biases. For instance, internet-based datasets can include unmoderated, biased, or hateful content. Academic papers tend to be more neutral, but they often lack the diversity of language found in social media or everyday conversations (Gallegos et al., 2024; KUNTZ and SILVA, 2023; Qu and Wang, 2024). Choosing which texts to include or exclude during the data collection phase can also lead to biases (Navigli et al., 2023). For instance, excluding certain groups or viewpoints, intentionally or unintentionally, can introduce selection bias, which can skew the model’s outputs toward the dominant narrative in the training data. Even with careful curation by LLM creators, biases can still be embedded in the model’s learning, as it is nearly impossible to fully eliminate inappropriate content given the vast scale of the training corpus.
2.1.3 Bias in Language Contexts
@@
Biases can also stem from the context of language. Human languages often contain ambiguities (including lexical ambiguity, syntax ambiguity, semantic ambiguity, and context-dependent meaning) due to the flexible nature of words, syntax, and context, which can lead to model biases. For example, gender-neutral pronouns may associated with one gender due to the patterns in the training dataset (Kotek et al., 2023; Dwivedi et al., 2023). Certain linguistic constructions, such as figures of speech and rhetoric, add another layer of complexity and ambiguity to human language by using words and expressions in non-literal or persuasive ways and hence introduce biases (de Sá et al., 2024).
@@ 3.1 Natural Language Understanding (NLU) Tasks
Common manifestations of bias in NLU tasks include:
  • Gender Bias: Models may incorrectly associate certain professions or roles with a specific gender, leading to errors in tasks like coreference resolution (Zhao et al., 2018a). For example, assuming a doctor is male and a nurse is female, regardless of context.
@@ 3.2 Natural Language Generation (NLG) Tasks
NLG tasks involve generating coherent and contextually relevant text based on input or instructions (Chang et al., 2024). Extrinsic bias in these tasks can cause models to produce biased language or reinforce stereotypes in generated text
Common manifestations of bias in NLG tasks include:
  • Gender Bias: Models might generate responses that align with gender stereotypes, such as using male pronouns for leaders and female pronouns for nurturing roles (Zhao et al., 2018a).
@@
These biases affect tasks like question answering, sentence completion, conversational agents, recommender systems, machine translation, and summarization. Recurrent issues include reinforcing harmful stereotypes, underrepresenting minority cultures, and providing biased recommendations or translations.
In general, extrinsic biases in LLMs impact both understanding and generation tasks, leading to unfair or discriminatory outcomes. Addressing these biases is essential to ensure models perform fairly and accurately in real-world applications. An overview of these biases in NLU tasks is provided in Table 1, and in NLG tasks in Table 2. Detailed examples can be found in Appendix A.
Table 1: Overview of extrinsic biases in natural language understanding (NLU) tasks.
@@
1) Stereotypical Occupation Associations: Mislinking pronouns based on gender stereotypes in professions, e.g., assuming doctors are male (Zhao et al., 2018a)
2) Gendered Pronoun Resolution: Incorrectly resolving pronouns for gender-neutral names, reflecting gender biases (Rudinger et al., 2018)
1) Assumptions About Technological Proficiency: Attributing technological skill to certain ages when resolving pronouns (Hovy and Yang, 2021)
2) Bias in Linking Pronouns to Age-Related Roles: Misassigning actions to younger or older individuals based on stereotypes (Garg et al., 2018)
1) Regional Variants and Pronoun Use: Difficulty with pronoun resolution in pro-drop languages due to training on non-pro-drop data (Hovy and Yang, 2021)
2) Cultural Context in Family Roles: Misresolving pronouns based on cultural stereotypes about family responsibilities (Rudinger et al., 2018)
@@
1) Gendered Language and Pronoun Resolution: Lower similarity scores for sentences differing only by gendered pronouns (Zhao et al., 2017)
2) Gender-Stereotyped Professions: Overestimating similarity when sentences align with gender stereotypes (Bolukbasi et al., 2016)
1) Age-Based Expectations in Language Use: Bias in similarity assessments due to age-related stereotypes (Caliskan et al., 2017)
2) Age Stereotypes in Sentiment and Perception: Overestimating similarity based on stereotypes about age and wisdom (Diaz et al., 2016)
1) Cultural Idioms and Expressions: Underestimating similarity between culturally equivalent idioms (Vulic and Moens, 2013)
2) Regional Dialects and Variants: Misjudging similarity due to differences in regional vocabulary (Tan and Bond, 2011)
@@
1) Stereotypical Gender Roles in Professions: Incorrect inferences based on gender stereotypes in job roles (Rudinger et al., 2018)
2) Gendered Language and Assumptions: Bias in inference due to gender assumptions with neutral names (Zhao et al., 2017)
1) Bias Toward Younger Individuals in Dynamic Roles: Overemphasizing youth in roles like entrepreneurship (Caliskan et al., 2017)
2) Age-Related Stereotypes in Professional Contexts: Assuming leadership roles are held by older individuals (Zhao et al., 2018a)
1) Cultural Norms and Social Roles: Inferences influenced by cultural stereotypes about professions (Rudinger et al., 2018)
2) Regional Political and Historical Contexts: Difficulty with inferences requiring diverse historical knowledge (Chen et al., 2016)
@@
1) Gendered Language in Job Title Classification: Associating certain professions with specific genders (Bolukbasi et al., 2016)
2) Gendered Pronouns and Name Classification: Misclassifying names or pronouns based on gender stereotypes (Zhao et al., 2017)
1) Sentiment Classification Bias Based on Age: Misclassifying sentiment due to age-related stereotypes (Díaz et al., 2018)
2) Bias in Job Application Classification: Discriminating against older applicants in job screening (De-Arteaga et al., 2019)
1) Misclassification Due to Cultural Language Variants: Incorrectly classifying non-standard dialects as negative (Sap et al., 2019a)
2) Regional Bias in Political Text Classification: Bias toward dominant regional political ideologies (Bender et al., 2021)
@@
1) Stereotypes in Gendered Activities: Assuming gender based on activities, e.g., cooking linked to females (Webster et al., 2018)
2) Assumptions About Gender Roles in Family Settings: Inferring roles based on traditional gender norms (Caliskan et al., 2017)
1) Bias in Health-Related Content: Associating certain health issues with age stereotypes (Bolukbasi et al., 2016)
2) Misinterpretation of Age-Related Roles: Incorrectly assuming roles based on age (Bender and Friedman, 2018)
1) Cultural Context Misinterpretation: Misunderstanding culturally specific practices in texts (Bender and Friedman, 2018)
2) Regional Language Varieties and Dialects: Struggling with comprehension of non-standard dialects (Blodgett et al., 2020)
@@
1) Bias in Sentiment Toward Gendered Products or Topics: Skewed sentiment analysis for gender-specific items (Zhao et al., 2017)
2) Sentiment Analysis in Gendered Contexts: Misclassifying sentiment in discussions challenging gender norms (Misiunas and Keyser, 2019)
1) Sentiment Analysis on Age-Related Topics: Assuming negative sentiment in texts by older individuals (Burnap and Williams, 2015)
2) Stereotyping Language Use by Older Adults: Misinterpreting sentiment due to age-related language stereotypes (Hovy and Søgaard, 2015)
1) Cultural Bias in Sentiment Toward Social Norms: Misclassifying sentiment across different cultural norms (Blodgett and O’Connor, 2017)
2) Sentiment in Multilingual Contexts: Incorrect sentiment assessment in code-switching texts (Hovy and Søgaard, 2015)
Table 2: Overview of extrinsic biases in natural language generation (NLG) tasks.
@@
1) Gender Bias in Answer Generation: When asked, “What should a good leader do?”, the model might use stereotypically male attributes, implying leadership qualities are inherently male (Zhao et al., 2018a)
2) Bias in Answering Ambiguous Gender Questions: For questions without specified gender, the model might default to male pronouns, reinforcing gender assumptions (Webster et al., 2018)
1) Stereotypical Answers About Aging: When asked about activities for elderly people, the model might focus on sedentary activities, neglecting active pursuits (Caliskan et al., 2017)
2) Negative Bias Toward Youth: Suggesting that young people lack experience for roles like managing a company (Bolukbasi et al., 2016)
1) Bias in Answering Culturally Specific Questions: Providing answers based on regional popularity rather than global knowledge, e.g., stating “American football” as the most popular sport (Gardner et al., 2018)
2) Language and Regional Bias in Answer Accuracy: Better performance on questions related to well-represented regions and languages (Kwiatkowski et al., 2019)
@@
1) Gender Bias in Descriptions of Physical Appearance: Completing sentences in ways that reinforce stereotypes about women’s and men’s concerns (Caliskan et al., 2017)
2) Gender Bias in Personal Attributes: Associating women with “emotional” and men with “strong” in sentence completions (Zhao et al., 2017)
1) Activity and Lifestyle Assumptions: Completing sentences with stereotypical activities based on age (Garg et al., 2018)
2) Learning and Education Stereotypes: Assuming older adults are “catching up” on education (Sun et al., 2019)
1) Cultural Stereotyping in Sentence Completion: Overemphasizing specific aspects of a culture, e.g., “In Japan, people often eat sushi” (Hendricks et al., 2018)
2) Regional Bias in Place-Based Completions: Reinforcing stereotypes about regions, e.g., “In Africa, many people live in villages” (Birhane and Prabhu, 2021)
@@
Generating coherent dialogue and maintaining conversation context (Zhang et al., 2019b)
1) Stereotypical Responses Based on Gendered Prompts: Reinforcing gender stereotypes in descriptions, e.g., nurses as female and caring (Bolukbasi et al., 2016)
2) Bias in Gendered Interactions: Responding differently based on assumed user gender (Hitsch et al., 2010)
1) Assumptions About Technological Savvy: Assuming tech-savvy users are younger and providing simplistic explanations to older users (Hovy and Søgaard, 2015)
2) Bias in Addressing Age-related Concerns: Discouraging older individuals from pursuing new careers (Wagner et al., 2016)
1) Culturally Inappropriate Responses: Failing to understand cultural norms in responses (Hovy and Spruit, 2016)
2) Bias in Handling Regional Topics: Focusing on negative topics for certain regions (Gebru et al., 2021)
@@
Suggesting personalized content or items based on user data (Zhang et al., 2019a)
1) Product Recommendations Based on Gender Stereotypes: Suggesting products reinforcing traditional gender roles (Ekstrand et al., 2018)
2) Career and Education Recommendations: Suggesting STEM careers more to men and arts to women (Chen et al., 2019)
1) Age-Related Product Recommendations: Recommending products based on age stereotypes (Biega et al., 2018)
2) Media and Entertainment Recommendations: Assuming preferences based on age, limiting content diversity (Burke et al., 2018)
1) Regional Bias in News and Information Recommendations: Under-representing news from minority regions (Karimi et al., 2018)
2) Bias in Language and Cultural Content Recommendations: Prioritizing content in dominant languages (Lakew et al., 2018)
@@
1) Gendered Language Mismatch: Introducing gender bias when translating from gender-neutral to gendered languages (Prates et al., 2020)
2) Gender Stereotyping in Occupational Translations: Assigning gendered pronouns based on stereotypes (Neri et al., 2020)
1) Bias in Addressing Older Adults: Translations that condescend to older adults (Caliskan et al., 2017)
2) Translation of Age-Related Idioms: Reinforcing negative stereotypes in translations (Hovy and Søgaard, 2015)
1) Cultural Nuance Loss: Mistranslating idioms and expressions without cultural context (Vanmassenhove et al., 2019)
2) Bias Toward Dominant Cultures: Favoring translations aligning with Western norms (Sennrich, 2016)
@@
1) Differential Emphasis on Roles: Emphasizing traditional gender roles in summaries (Bender and Friedman, 2018)
2) Selective Emphasis on Gendered Information: Overemphasizing gender-specific details not central to the story (Otterbacher et al., 2017)
1) Bias in Summarizing Content for Different Age Groups: Simplifying content in a condescending way (Obermeyer et al., 2019)
2) Omission of Contributions Based on Age: Highlighting contributions of younger people over older individuals (Bender and Friedman, 2018)
1) Omission of Culturally Significant Details: Omitting culturally important information in summaries (Shen et al., 2017)
2) Bias Toward Western Narratives: Prioritizing Western perspectives in global news summaries (Perez-Beltrachini and Lapata, 2022)

### v2

@@ 2.2 Bias in Data Collection Methods
The intrinsic biases can also stem from the methods that are used to collect the training corpus. A large training corpus of LLMs often comprises sub-datasets that are collected from different sources, including web scraping, social media and forums, books and literature, code repositories, scientific and academic papers, legal and finance documents, dialogue and transcriptions, government and public records, custom and proprietary databases, and model-generated datasets. These different types of sources have their own biases. For instance, internet-based datasets can include unmoderated, biased, or hateful content. Academic papers tend to be more neutral, but they often lack the diversity of language found in social media or everyday conversations (Gallegos et al., 2024; KUNTZ and SILVA, 2023; Qu and Wang, 2024). Choosing which texts to include or exclude during the data collection phase can also lead to biases (Navigli et al., 2023). For instance, excluding certain groups or viewpoints, intentionally or unintentionally, can introduce selection bias, which can skew the model’s outputs toward the dominant narrative in the training data. Even with careful curation by LLM creators, biases can still be embedded in the model’s learning, as it is nearly impossible to fully eliminate inappropriate content given the vast scale of the training corpus.
2.3 Bias in Language Contexts
@@
Biases can also stem from the context of language. Human languages often contain ambiguities (including lexical ambiguity, syntax ambiguity, semantic ambiguity, and context-dependent meaning) due to the flexible nature of words, syntax, and context, which can lead to model biases. For example, gender-neutral pronouns may be associated with one gender due to the patterns in the training dataset (Kotek et al., 2023; Dwivedi et al., 2023). Certain linguistic constructions, such as figures of speech and rhetoric, add another layer of complexity and ambiguity to human language by using words and expressions in non-literal or persuasive ways and hence introduce biases (de Sá et al., 2024).
2.4 Bias in Tokenization and Word Embeddings
Tokenization and the use of word embeddings can further introduce biases into LMs. Some tokenization techniques, such as WordPiece or Byte-Pair Encoding (BPE), may disproportionately split less frequently occurring words into smaller units. This splitting policy can result in fragmented representations of under-represented entities, names, or terminology, particularly affecting minority languages or groups. For instance, Petrov et al. show that biases can arise at the tokenization stage and that the same content can require substantially different token lengths across languages Petrov et al. (2023). Ahia et al. demonstrate large cross-lingual token-count disparities with practical fairness consequences Ahia et al. (2023). Ovalle et al. Ovalle et al. (2024) discover that Byte-Pair Encoding (BPE) tokenization, the tokenizer powering many popular LLMs, influences LLM misgendering significantly. Similarly, biases can also arise from word embedding, where many language models are initialized with pretrained word embeddings Bolukbasi et al. (2016); Caliskan et al. (2017); Garg et al. (2018); Nadeem et al. (2021); Rakshit et al. (2025).
@@ 3.1 Natural Language Understanding (NLU) Tasks
Common manifestations of bias in NLU tasks include the following:
  • Gender bias: Models may incorrectly associate certain professions or roles with a specific gender, leading to errors in tasks like coreference resolution (Zhao et al., 2018). For example, assuming a doctor is male and a nurse is female, regardless of context.
@@ 3.2 Natural Language Generation (NLG) Tasks
NLG tasks involve generating coherent and contextually relevant text based on input or instructions (Chang et al., 2024). Extrinsic bias in these tasks can cause models to produce biased language or reinforce stereotypes in generated text.
Common manifestations of bias in NLG tasks include the following:
  • Gender bias: Models might generate responses that align with gender stereotypes, such as using male pronouns for leaders and female pronouns for nurturing roles (Zhao et al., 2018).
@@
These biases affect tasks like question answering, sentence completion, conversational agents, recommender systems, machine translation, and summarization. Recurrent issues include reinforcing harmful stereotypes, under-representing minority cultures, and providing biased recommendations or translations.
In general, extrinsic biases in LLMs impact both understanding and generation tasks, leading to unfair or discriminatory outcomes. Addressing these biases is essential to ensure models perform fairly and accurately in real-world applications. An overview of these biases in NLU tasks is provided in Table LABEL:tab:nlu-biases, and in NLG tasks in Table LABEL:tab:nlg-biases. Detailed examples can be found in Appendix A.
Table 1: Overview of extrinsic biases in natural language understanding (NLU) tasks. Each row corresponds to an NLU task, and each column summarizes representative examples of gender, age, and cultural or regional bias. The examples highlight typical patterns rather than exhaustive cases. This table can be used to compare how different types of bias manifest across tasks.
@@
(1) Stereotypical occupation associations: Mislinking pronouns based on gender stereotypes in professions, e.g., doctor → male (Zhao et al., 2018)
(2) Gendered pronoun resolution: Incorrectly resolving pronouns for gender-neutral names, reflecting underlying gender biases (Rudinger et al., 2018)
(1) Assumptions about technological proficiency: Attributing technological skill to certain ages during pronoun resolution (Hovy and Yang, 2021)
(2) Bias in linking pronouns to age-related roles: Misassigning actions to younger or older individuals based on stereotypes (Garg et al., 2018)
(1) Regional variants and pronoun use: Difficulty resolving pronouns in pro-drop languages due to training on non-pro-drop data (Hovy and Yang, 2021)
(2) Cultural context in family roles: Misresolving pronouns based on cultural stereotypes about family responsibilities (Rudinger et al., 2018)
@@
(1) Gendered language and pronoun resolution: Lower similarity scores for sentences differing only by gendered pronouns (Zhao et al., 2017)
(2) Gender-stereotyped professions: Overestimating similarity when sentences align with gender stereotypes (Bolukbasi et al., 2016)
(1) Age-based expectations in language use: Bias in similarity assessments due to age-related stereotypes (Caliskan et al., 2017)
(2) Age stereotypes in sentiment and perception: Overestimating similarity based on stereotypes about age and wisdom (Diaz et al., 2016)
(1) Cultural idioms and expressions: Underestimating similarity between culturally equivalent idioms (Vulic and Moens, 2013)
(2) Regional dialects and variants: Misjudging similarity due to differences in regional vocabulary (Tan and Bond, 2011)
@@
(1) Stereotypical gender roles in professions: Incorrect inferences based on gender stereotypes in job roles (Rudinger et al., 2018)
(2) Gendered language and assumptions: Bias in inference due to gender assumptions with neutral names (Zhao et al., 2017)
(1) Bias toward younger individuals in dynamic roles: Overemphasizing youth in roles like entrepreneurship (Caliskan et al., 2017)
(2) Age-related stereotypes in professional contexts: Assuming leadership roles are held by older individuals (Zhao et al., 2018)
(1) Cultural norms and social roles: Inferences influenced by cultural stereotypes about professions (Rudinger et al., 2018)
(2) Regional political and historical contexts: Difficulty with inferences requiring diverse historical knowledge (Chen et al., 2016)
@@
(1) Gendered language in job title classification: Associating certain professions with specific genders (Bolukbasi et al., 2016)
(2) Gendered pronouns and name classification: Misclassifying names or pronouns based on gender stereotypes (Zhao et al., 2017)
(1) Sentiment classification bias based on age: Misclassifying sentiment due to age-related stereotypes (Díaz et al., 2018)
(2) Bias in job application classification: Discriminating against older applicants in job screening (De-Arteaga et al., 2019)
(1) Misclassification due to cultural language variants: Incorrectly classifying non-standard dialects as negative (Sap et al., 2019)
(2) Regional bias in political text classification: Bias toward dominant regional political ideologies (Bender et al., 2021)
@@
(1) Stereotypes in gendered activities: Assuming gender based on activities, e.g., cooking → female (Webster et al., 2018)
(2) Assumptions about gender roles in family settings: Inferring roles based on traditional gender norms (Caliskan et al., 2017)
(1) Bias in health-related content: Associating certain health issues with age stereotypes (Bolukbasi et al., 2016)
(2) Misinterpretation of age-related roles: Incorrectly assuming roles based on age (Bender and Friedman, 2018)
(1) Cultural context misinterpretation: Misunderstanding culturally specific practices in texts (Bender and Friedman, 2018)
(2) Regional language varieties and dialects: Struggling with comprehension of non-standard dialects (Blodgett et al., 2020)
@@
(1) Bias in sentiment toward gendered products or topics: Skewed sentiment analysis for gender-specific items (Zhao et al., 2017)
(2) Sentiment analysis in gendered contexts: Misclassifying sentiment in discussions challenging gender norms (Misiunas and Keyser, 2019)
(1) Sentiment analysis on age-related topics: Assuming negative sentiment in texts by older individuals (Burnap and Williams, 2015)
(2) Stereotyping language use by older adults: Misinterpreting sentiment due to age-related language stereotypes (Hovy and Søgaard, 2015)
(1) Cultural bias in sentiment toward social norms: Misclassifying sentiment across different cultural norms (Blodgett and O’Connor, 2017)
(2) Sentiment in multilingual contexts: Incorrect sentiment assessment in code-switching texts (Hovy and Søgaard, 2015)
Table 2: Overview of extrinsic biases in natural language generation (NLG) tasks. Each row corresponds to an NLG task, while columns summarize representative examples of gender, age, and cultural or regional bias. These examples illustrate common bias patterns and are not exhaustive. Table can be used to compare how different types of bias manifest across tasks.
@@
(1) Gender bias in answer generation: When asked, “What should a good leader do?”, the model might use stereotypically male attributes, implying leadership qualities are inherently male (Zhao et al., 2018)
(2) Bias in answering ambiguous gender questions: For questions without specified gender, the model might default to male pronouns, reinforcing gender assumptions (Webster et al., 2018)
(1) Stereotypical answers about aging: When asked about activities for elderly people, the model might focus on sedentary activities, neglecting active pursuits (Caliskan et al., 2017)
(2) Negative bias toward youth: Suggesting that young people lack experience for roles like managing a company (Bolukbasi et al., 2016)
(1) Bias in answering culturally specific questions: Providing answers based on regional popularity rather than global knowledge, e.g., stating “American football” as the most popular sport (Gardner et al., 2018)
(2) Language and regional bias in answer accuracy: Better performance on questions related to well-represented regions and languages (Kwiatkowski et al., 2019)
@@
(1) Gender bias in descriptions of physical appearance: Completing sentences in ways that reinforce stereotypes about women’s and men’s concerns (Caliskan et al., 2017)
(2) Gender bias in personal attributes: Associating women with “emotional” and men with “strong” in sentence completions (Zhao et al., 2017)
(1) Activity and lifestyle assumptions: Completing sentences with stereotypical activities based on age (Garg et al., 2018)
(2) Learning and education stereotypes: Assuming older adults are “catching up” on education (Sun et al., 2019)
(1) Cultural stereotyping in sentence completion: Overemphasizing specific aspects of a culture, e.g., “In Japan, people often eat sushi” (Hendricks et al., 2018)
(2) Regional bias in place-based completions: Reinforcing stereotypes about regions, e.g., “In Africa, many people live in villages” (Birhane and Prabhu, 2021)
@@
Generating coherent dialogue and maintaining conversation context (Zhang et al., 2019)
(1) Stereotypical responses based on gendered prompts: Reinforcing gender stereotypes in descriptions, e.g., nurses as female and caring (Bolukbasi et al., 2016)
(2) Bias in gendered interactions: Responding differently based on assumed user gender (Hitsch et al., 2010)
(1) Assumptions about being tech-savvy: Assuming tech-savvy users are younger and providing simplistic explanations to older users (Hovy and Søgaard, 2015)
(2) Bias in addressing age-related concerns: Discouraging older individuals from pursuing new careers (Wagner et al., 2016)
(1) Culturally inappropriate responses: Failing to understand cultural norms in responses (Hovy and Spruit, 2016)
(2) Bias in handling regional topics: Focusing on negative topics for certain regions (Gebru et al., 2021)
@@
Suggesting personalized content or items based on user data (Zhang et al., 2019)
(1) Product recommendations based on gender stereotypes: Suggesting products reinforcing traditional gender roles (Ekstrand et al., 2018)
(2) Career and education recommendations: Suggesting STEM careers more to men and arts to women (Chen et al., 2019)
(1) Age-related product recommendations: Recommending products based on age stereotypes (Biega et al., 2018)
(2) Media and entertainment recommendations: Assuming preferences based on age, limiting content diversity (Burke et al., 2018)
(1) Regional bias in news and information recommendations: Under-representing news from minority regions (Karimi et al., 2018)
(2) Bias in language and cultural content recommendations: Prioritizing content in dominant languages (Lakew et al., 2018)
@@
(1) Gendered language mismatch: Introducing gender bias when translating from gender-neutral to gendered languages (Prates et al., 2020)
(2) Gender stereotyping in occupational translations: Assigning gendered pronouns based on stereotypes (Neri et al., 2020)
(1) Bias in addressing older adults: Translations that condescend to older adults (Caliskan et al., 2017)
(2) Translation of age-related idioms: Reinforcing negative stereotypes in translations (Hovy and Søgaard, 2015)
(1) Cultural nuance loss: Mistranslating idioms and expressions without cultural context (Vanmassenhove et al., 2019)
(2) Bias toward dominant cultures: Favoring translations aligning with dominant (e.g., Western) norms (Sennrich, 2016)
@@
(1) Differential emphasis on roles: Emphasizing traditional gender roles in summaries (Bender and Friedman, 2018)
(2) Selective emphasis on gendered information: Overemphasizing gender-specific details not central to the story (Otterbacher et al., 2017)
(1) Bias in summarizing content for different age groups: Simplifying content in a condescending way (Obermeyer et al., 2019)
(2) Omission of contributions based on age: Highlighting contributions of younger people over older individuals (Bender and Friedman, 2018)
(1) Omission of culturally significant details: Omitting culturally important information in summaries (Shen et al., 2017)
(2) Bias toward western narratives: Prioritizing Western perspectives in global news summaries (Perez-Beltrachini and Lapata, 2022)